OpenStudy (anonymous):

A right triangle has sides measuring 6 units, 8 units, and 10 units. Which of the following shows the Pythagorean relationship of the areas of the squares drawn along each side? A. 12 + 16 = 28 B. 18 + 24 = 42 C. 24 + 76 = 100 D. 36 + 64 = 100 I do not want the answer or it explained to me yet. I wanna know what the question is ASKING. :) I don't get what it's asking...

OpenStudy (anonymous):

that looked right to me...

OpenStudy (xguardians):

Sorry, its asking for what the Pythagorean Theorem will look like based off those numbers.

OpenStudy (anonymous):

\(6^2+8^2=10^2\) is what they are looking for

OpenStudy (xguardians):

@satellite73 "I do not want the answer or it explained to me yet. I wanna know what the question is ASKING. :) I don't get what it's asking..."

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

Don't give me the answer please... These are study questions for a major test and I'm trying to get through them. :) I HAVE to know this to move onto the next grade.

OpenStudy (anonymous):

it is asking for the areas of the squares built on the sides of the triangle the area of a square is the square of the length of the sides

OpenStudy (anonymous):

It's just saying which one shows 6^2+8^2=10^2

OpenStudy (anonymous):

which is D.

OpenStudy (anonymous):

so if one side has length \(6\) the area is \(6^2\)

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

the Pythagorean therom is a^2 + b^2 = c^2

OpenStudy (anonymous):

@CrazyCountryGirl do you understand ?

OpenStudy (texaschic101):

a^2 + b^2 = c^2 the longest side is the hypotenuse...and it is c the other sides are a and b and they measure 6 and 8 now just plug them into the formula